Shift Summary - DAY - Sept 20, 1999

Douglas A. Jensen (djensen@fnal.gov)
Mon, 20 Sep 1999 16:26:24 -0500 (CDT)

Shift Summary for DAY shift, 20/Sept/99

On Shift: Ramberg, Jensen

Hours of beam available at all intensities: 8:00/8:30
Hours of beam available at nominal intensity: 8:00
Hours of E799 data (special runs not included): 6:30/ 8:30
Hours of special runs (laser scans, muon runs...): 1:00

Some beam off. A bunch of restarts of the DAQ, plugging in of
DAQ board, special runs.

========================================================================
Runs ended this shift: 14643,47,48,49,50,51

Tapes brought to FCC this shift: none.

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Took a block of E799 kaon data. Special runs for Hyperon Trigger
Studies. Had some difficulties restarting DART, turned out to be
a loose board in the KUMQUAT rack.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14643 ~62 ~07:00 New shift on board. Running with low (machine
limited) intensity
77 no beam
86 beam back
91 07:45 ~8E12, up from ~4E12
08:50 mounted UPX001 on ktev3 as per messages from split pgm
09:00 mounted UPF001 on ktev2 as per message.
137 09:55 missed about 6 spills.
190 09:58 Note that spill counting etc on DA is frozen, so
spill no, time etc not updating.
>> trying to sort out L2 STT accepts, random, prescaled....
214 10:30 mucking with stt random accept cables
221 10:40 end the run. Last few spills mucking with
trigger and prescale etc.

Need to do a DART fresh start to get display working
again. As a result of tests, Bob T says we can
change prescale on Trigger 10 to 1 ( from 2 )
Jason H implements same.
Upon start of run, kumquots hung. reset, DART rest.
11:15 kumquots still hung, after reload, fresh start,
Jason L paged. He finds caontrol card unplugged !

14647 1 11:57 Start run
3 12 UR quadrant disappears from on-line pictures.
lost 1st 5 spills of the run for this one.
Note Trig 10
14647 100 14:15 End run


14648 1 14:30 Start - Hyperon Trigger Studies
>=1 HCC, HAveto
20


14649 1 Hyperon Trigger studies,
>=2 HCC, HAveto
15:02 UPM002 mounted on tps3d2 on KTeV1 Split mount
20 15:20 end run, beam off for about 8 spills

15:30 first of replacement crew arrives.

14650 1 15:34 2HCX ( 2 in each x chamber ), HAveto, HC <= 1
a few KFRIENDS and reload to get going
16 15:53 end run

14651 1 ~16:00 2HCX, HA not in veto, HC <=2
13 16:15 end the run

We will then go to HA not in veto, HC <= 1; 2HCX not in
while continuing the run at low intensity.

=========================================================================
Follow-up from previous shift:
CV and KUMQUOT red voltage monitors - known problems, may ignore.
Picked up a couple of toner carts. Installed one.
=========================================================================
Follow-up on next shift:

1)Clean out the KTeV refrigerator, before the whole collaboration gets
sick. < I didn't do this (daj) - people should clean up their own messes!

2)Look into debugging trigger 16 which is still not working.

3) Question for tracking experts - why in the plots - do we see about
3% of events with >= 40 vertices ! Seems strange. ????

4) The Xerox machine is still not working. Some discussion of how
it could be that this is such a difficult problem......

Douglas Jensen DJensen@fnal.gov x8194